OpenStudy (anonymous):

is this right?: (g+4)-3g = 1+g (4+g)-3g=1+g 4+g-3g=1+g 4+(-3g)=1+g 4+1= -3g+g 5=-2g -- --- (divide) -2 -2 g=-2.5 I don't think it's right. help please

OpenStudy (anonymous):

It should be... (g+4)-3g = 1+g -2g + 4 = 1 + g 4 = 1 + 3g 3 = 3g 1 = g

OpenStudy (anonymous):

Oh! oops, I accidentally put g+-3g as -3g. Hehe... Thank you!!

OpenStudy (anonymous):

youve got it right i think,